OpenStudy (bleuspectre):

PLEASE HELP Find the center vertices and foci of each ellipse x2/16 + y2/25 = 1? A.center (0,0), vertices (0,-5), (0,5), Foci (0,-3) (0,3) B.center (0,0) vertices (-5,0),(5,0) Foci (-3,0), (3,0)

OpenStudy (mathmale):

Happy to help, but I do have a couple of up-front requests. First, please post only one problem at a time. Second, think about the problem, look for examples to follow, share your thoughts and any work you've done, and then ask questions here on OpenStudy. x2/16 + y2/25 = 1 represents the equation of an ellipse which is longer vertically than it is wide. Representing the square of x by x^2 (not x2) and the square of y by y^2, this equation rewritten is x^2 y^2 ---- + ---- = 1 4^2 5^2 or \[\frac{ x^2 }{ 4^2 }+\frac{ y^2 }{ 5^2 }=1\]

OpenStudy (mathmale):

At this point I'm going to ask you to look up in your textbook (if you have one) or onlilne the topic "Equation of an ellipse." Most explanations will include drawings which will show the dimensions a, b and c. 2a is the length of the major (longer) axis of the ellipse, whereas 2b is that of the minor (shorter) axis. You'll need to find (through your research) an equation that relates a, b and c for an ellipse. This is a lot to digest! Ask specific questions when you're ready.

OpenStudy (bleuspectre):

I did ask a specific question.

OpenStudy (bleuspectre):

This was the question i got asked for homework. I dont understand it.

OpenStudy (mathmale):

My suggestion is that you do some preliminary work before asking for help. Can you, for example, define "focus" and "vertex" in the context of an ellipse? Can you define what a, b and c represent? I do expect to see some basic research on your part, or specific questions about the problem solving process that applies here, before giving help.

OpenStudy (mathmale):

It clearly shows how a, b and c are related to one another and what they measure.

OpenStudy (mathmale):

Take a look at the equation for an ellipse that y ou've posted. What is the value of a? of b? of c? Does the ellipse appear to be vertical or horizontal, and how do you know? What are the coordinates of the center? of the vertices? of the foci?